# Language-Driven Policy Distillation for Cooperative Driving in Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Index Terms:
Cooperative Decision-Making, large language model, multi-agent reinforcement learning
@@ I Introduction
Recent advancements in Connected and Autonomous Vehicle (CAV) technology have enabled vehicles to communicate and collaborate, revolutionizing the transportation industry by enhancing overall traffic management[1, 2]. Cooperative driving technology is a critical component of CAV deployment, which enables them to make efficient decisions in dynamic environments, thereby enhancing safety and traffic flow. However, despite the significant potential of cooperative decision-making technology, current methods still face numerous challenges in terms of scenario applicability, decision-making efficiency, and safety, making large-scale real-world deployment a distant goal[3].
@@
Recent advancements in deep learning have made learning-based approaches, such as Multi-Agent Reinforcement Learning (MARL), effective for cooperative decision-making[4]. In MARL, each agent explores strategies to maximize its own reward, thereby identifying optimal decision-making policies of CAV, which has been employed by many researchers[5, 6, 7]. Despite the notable advantages of MARL, challenges remain, such as difficulties in algorithm convergence, low exploration efficiency, and high sampling costs[7]. On the other hand, recent advancements in Large Language Models (LLMs) have demonstrated impressive zero-shot generalization and complex logical reasoning capabilities across various decision-making tasks[8]. These capabilities are particularly well-suited for assessing complex situations in human traffic environments and achieving coordinated control between vehicles and traffic systems. A few studies have explored the integration of LLMs into traffic and vehicle systems to enhance the safety and efficiency of transportation networks[9, 10]. However, in real-time, dynamic traffic scenarios, existing LLMs struggle to enable CAVs to make rapid and efficient decisions. Additionally, the high deployment costs hinder large-scale implementation, significantly limiting the potential of LLMs[11].
@@
To address these challenges, we propose a novel approach that leverages the reasoning efficiency of MARL-based agents alongside the extensive world knowledge of LLMs. This approach involves designing a teacher-student policy distillation framework to facilitate the efficient transfer of LLMs’ world knowledge to MARL agents. We introduce LDPD, a language-driven policy distillation and learning framework designed to guide MARL exploration. As illustrated in Fig.2, this framework comprises a teacher agent based on LLM and multiple student agents (CAVs) based on MARL. During the training phase, the teacher agent receives observation information from the Students, leveraging its powerful zero-shot and reasoning capabilities to make expert-level decisions. It then guides the smaller student agents through its actions, enabling them to perform efficient sampling and policy learning. After the learning phase, these Student policies will function as the controllers for CAVs, making decisions independently of the teacher during deployment.
For teacher agent, the Observation module preprocesses and enhances student observations, which are then sent to the LLM-based Planner for decision-making. The Planner integrates the data using agent tools, and expert-level outputs are generated. These expert decisions are used to train the student policies. The student agents are trained in a decentralized MARL framework, using the Actor-Critic method to update their action policies.
@@ II Related Works
Cooperation Decision-Making of CAVs. In complex mixed-autonomy traffic environments, cooperative decision-making technologies can effectively improve the efficiency and safety of the CAV system. Numerous methods have been proposed to address vehicle coordination, including rule-based methods[12], optimized-based methods[13, 14], and learning-based methods[15, 16]. Among them, learning-based methods are highly effective in simulating vehicle interaction dynamics and reasoning, while their lack of interpretability and limited generalization capabilities hinder their application in safety-critical scenarios.
Large Language Model for Decision-making. Large Language Models (LLMs), known for their human-like reasoning and commonsense generalization, have recently demonstrated potential beyond traditional natural language processing tasks [17, 18]. In recent years, LLMs have been utilized as employed as end-to-end agents capable of autonomous decision-making. For cooperative driving tasks, Hu et al. proposed AgentsCoDriver[10], a LLM-driven lifelong learning framework to empower CAV decision-making. Meanwhile, CoDrivingLLM, proposed by Fang et al.[9], is designed to adapt different levels of Cooperative Driving Automation, which could effectively address the delimma of cooperative scenarios. While LLMs have demonstrated remarkable potential, significant challenges remain. Their high computational demands and the complexity of real-time decision-making introduce latency issues, which limit their application in time-critical environments.
@@ III Problem Formulation
The cooperative decision-making problem for multiple CAVs can be modeled as a partially observable Markov decision process (POMDP)[22]. We define the POMDP using the tuple $\mathcal{M}_{\mathcal{G}}=(\mathcal{V},S,[\mathcal{O}_{i}],[\mathcal{A}_{i}],\mathcal{P},[r_{i}])$, where $\mathcal{V}$ represents the finite set of all controlled agents (CAVs), and $S$ denotes the state space encompassing all agents. $\mathcal{O}_{i}$ represents the observation space for each agent $i\in\mathcal{V}$, $\mathcal{A}_{i}$ denotes the action space, and $r_{i}$ is the reward associated with CAV $i$. The transition distribution is represented by $\mathcal{P}$.
At any given time, each agent $i$ receives an individual observation $o_{i}:\mathcal{S}\to\mathcal{O}_{i}$ and selects an action $a_{i}\in\mathcal{A}_{i}$ based on a policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$. The agent then transitions to a new state $s_{i}^{\prime}$ with a probability given by $\mathcal{P}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{N}\to\mathcal{S}$.
III-1 Observation Space
@@
where $x_{k}$ ,$y_{k}$ $v^{x}_{k}$ and $v^{y}_{k}$ are the longitudinal and lateral positions and speeds, respectively. $\phi_{k}$ is the vehicle’s heading angle. The overall observation space of the system is the combined observation of all CAVs, i.e., $\mathcal{O}=\mathcal{O}_{1}\times\mathcal{O}_{2}\times\cdots\times\mathcal{O}_{|\mathcal{V}|}$.
III-2 Action Space
[…]
III-3 Reward function
The reward function has a great effect on the performance of agents. In this paper, aiming to make all the agents pass the merging area safely and efficiently, the reward of $i$th agent at the time step $t$ is defined as follows:
@@ IV Method
In this section, the whole LDPD framework is first overviewed. Then the teacher agent and student agent are introduced respectively.
@@ IV-A Framework Overview
The LDPD framework is designed to facilitate the learning and exploration process of multi-agent systems with distilled knowledge from LLM, especially in the complex cooperative driving scenarios. As shown in Fig.2, the LDPD framework comprises two main components: the teacher agent and the student agents. The teacher agent facilitates efficient policy learning for multiple student agents through its action demonstrations. The student agents, composed of multiple small policy networks, form an agent group where each network is responsible for controlling an individual CAV. Throughout this process, the student agents efficiently distill and apply the knowledge from the LLM, enabling rapid learning and adaptation.
@@ IV-B1 Observation Enhancer
Additionally, based on the state information of the CAV, we predict and infer the driving intentions of each CAV, including the desired target lane and intended driving behavior. At time $t$, the driving intention information for CAV $i$ is represented as $In_{i,t}=\{lane_{i,t},behavior_{i,t}\}$, where $lane_{i,t}$ and $behavior_{i,t}$ refer to the desired lane intention and the desired behavior intention, respectively.
@@ IV-B2 Core Planner
We use the conflict checking function as an example to introduce our tools: After the LLM obtains the semantic description of the scenario, it will call the conflict checking tool $conflict\_check()$ before initiating reasoning to detect potential safety risks.
@@
The conflict checking tool first retrieves all potential conflicts in the current environment and assesses the risk level based on the current states of the conflicting parties. Here, we use the time to conflict point (TTCP) difference as a surrogate metric for the risk level, calculated as follows:
@@
$A=LLM(Sce,conflict\_info)$ (6)
@@ IV-B3 Safety Checker
Firstly, the formula is designed to calculate the priority score of each CAV:
@@
With the priority scores of CAVs calculated above, at time step $t$, the safety checker first produces a priority list ${P}_{t}$, which consists of a list of priority scores and their corresponding ego vehicles. Then, the safety checker sequentially examines the future trajectory of the vehicles in the list to determine if there are any conflicts between vehicles.
We use the Intelligent Driver Model (IDM)[24] to predict the longitudinal acceleration of HVs, based on their current speed and distance headway, and we utilize the MOBIL lane change model[25] to predict the lateral behavior of HVs.
@@
where ${A}_{available}$ is a set of available actions at time step $k$ for the selected agent, and ${d}_{sm,k}$ is the safety margin at prediction time step $k$.
The safety margin is defined as follows:
@@ V-B Implementation details
All experiments were conducted on a computation platform with an Intel Xeon Silver 4214R CPU, NVIDIA GeForce RTX 3090 GPU×$\times$2, and 128G Memory.
@@ V-C1 Overall Performance
The training curves for all methods under different conditions and scenarios are shown in Fig.5. As observed, in all experimental scenarios, our method (LEMAA2C) consistently outperform the MARL baselines. Furthermore, they demonstrate significantly higher exploration efficiency in the early stages of training, which effectively boosts the training efficiency of the student agents. This strongly validates the effectiveness of our approach.
Among all MARL methods, MAA2C, as used in this paper, exhibits the strongest overall performance, delivering the best results across most traffic densities and complexity levels, particularly excelling in the most challenging hard mode. In comparison, MAPPO performs slightly worse but still achieves good results in specific scenarios, such as scenario 2-easy. MAACKTR shows the weakest performance and adaptability, with its disadvantages becoming more pronounced as the scenario complexity increases.
@@ V-C2 Safety Performance
The model’s overall performance is tested under various experimental conditions, as shown in Tab.I. It is evident that our method consistently achieves the lowest collision rate across all tests. We also evaluate the safety performance of teacher agent under different conditions and compare it with the student agents, as shown in Tab. II. In most scenarios, LAMAA2C outperforms LLM in the final performance, demonstrating that while effectively learning from LLM, our method retains strong exploration capabilities to further improve performance.
TABLE I: Performance across scenarios for various methods under different difficulty settings
@@
TABLE II: Performance of teacher and students under varying difficulty levels.
@@ V-C3 Efficiency Performance
The average speed of all CAVs, shown in Tab.I, is analyzed to assess the efficiency of cooperative driving across different methods. We observe that as the difficulty of tasks increases, the average speed of the CAVs decreases, which likely represents a compromise made to ensure safety over efficiency. However, compared to other baseline methods, such as MAPPO and MAACKTR, our approach not only maintains better safety performance but also ensures driving efficiency. This demonstrates that our method effectively balances driving speed with safety.
@@ V-C4 Cross-Validation
To evaluate the adaptability and generalization capabilities of the learned strategies, we test the models in task environments beyond their training conditions. For example, models trained under the hard mode are now tested on easy and medium mode tasks, with results shown in Tab.III. We observe that models trained in hard mode not only outperform the original medium mode models in terms of safety and efficiency but also excel in other scenarios. This demonstrates the exceptional adaptability of our model.
TABLE III: Performance of students by original and applied levels across scenarios
Original Level Applied Level Evaluation Reward Collision Rate Average Speed Average PET
Sce-1 Sce-2 Sce-1 Sce-2 Sce-1 Sce-2 Sce-1 Sce-2
HARD Simple 62.66 46.18 0.02 0.13 24.90 24.85 11.29 7.95
Medium 51.16 70.78 0.07 0.00 23.77 24.87 7.35 21.43
MEDIUM Simple 69.31 61.64 0.00 0.02 25.26 26.52 11.74 17.37
@@ VI Conclusions
To enhance exploration and learning efficiency in cooperative decision-making for CAVs, we introduced LDPD, a language-driven policy distillation framework to facilitate MARL agents learning. Within this framework, the language model, as the teacher, makes complex reasoning decisions based on observations from student agents and stores these decision-making experiences. Student agents learn the teacher’s prior knowledge through gradient updates in their policy networks. Extensive experiments demonstrate that distilled infromation from the teacher significantly enhances the exploration efficiency of student agents. Compared to baseline methods, there is a notable overall improvement in the agents’ performance, eventually surpassing the expert decision-making level of the teacher.
Our proposed instructional framework is versatile and compatible with any MARL algorithm. In the future, we plan to test our method with more MARL algorithms and in a broader range of decision-making scenarios to further enhance the generalizability and adaptability of our approach.

### v2

@@ I Introduction
Recent advancements in Connected and Autonomous Vehicle (CAV) technology have enabled vehicles to communicate and collaborate, revolutionizing the transportation industry by enhancing overall traffic management [1, 2, 3]. Cooperative driving technology is a critical component of CAV deployment, which enables them to make efficient decisions in dynamic environments, thereby enhancing safety and traffic flow. However, despite the significant potential of cooperative decision-making technology, current methods still face numerous challenges in terms of scenario applicability, decision-making efficiency, and safety, making large-scale real-world deployment a distant goal [4].
@@
The rapid development in deep learning has made learning-based approaches, such as Multi-Agent Reinforcement Learning (MARL), effective for cooperative decision-making [5]. In MARL, each agent explores strategies to maximize its own reward, thereby identifying optimal decision-making policies of CAV, which has been employed by many researchers [6, 7, 8]. Despite the notable advantages of MARL, challenges remain, such as difficulties in algorithm convergence, low exploration efficiency, and high sampling costs [8]. On the other hand, recent advancements in Large Language Models (LLMs) have demonstrated impressive zero-shot generalization and complex logical reasoning capabilities across various decision-making tasks [9]. These capabilities are particularly well-suited for assessing complex situations in human traffic environments and achieving coordinated control between vehicles and traffic systems. A few studies have explored the integration of LLMs into traffic and vehicle systems to enhance the safety and efficiency of transportation networks [10, 11]. However, in real-time, dynamic traffic scenarios, existing LLMs struggle to enable CAVs to make rapid and efficient decisions. Additionally, the high deployment costs hinder large-scale implementation, significantly limiting the potential of LLMs [12].
@@
To address these challenges, we propose a novel approach that leverages the reasoning efficiency of MARL-based agents alongside the extensive world knowledge of LLMs. This approach involves designing a teacher-student policy distillation framework to facilitate the efficient transfer of LLMs’ world knowledge to MARL agents. we present LDPD, a language-driven policy distillation framework designed to guide MARL exploration. The framework features a teacher agent powered by LLM and multiple student agents (CAVs) modeled with MARL, enabling effective knowledge transfer and collaborative learning. During the training phase, the teacher agent receives observation information from students, leveraging its powerful zero-shot and reasoning capabilities to make expert-level decisions. It then guides the smaller student agents through its actions, enabling them to perform efficient sampling and policy learning. After the learning phase, these Student policies will function as the controllers for CAVs, making decisions independently of the teacher during deployment. For teacher agent, the Observation module preprocesses and enhances student observations, which are then sent to the LLM-based Planner for decision-making. The Planner integrates the data using agent tools, and expert-level outputs are generated. These expert decisions are used to train the student policies. The student agents are trained in a decentralized MARL framework, using the Actor-Critic method to update their action policies.
@@ II Related Works
Cooperation Decision-Making of CAVs. In complex mixed-autonomy traffic environments, cooperative decision-making technologies can effectively improve the efficiency and safety of the CAV system. Numerous methods have been proposed to address vehicle coordination, including rule-based methods [13], optimized-based methods [14], and learning-based methods [15]. Among them, learning-based methods are highly effective in simulating vehicle interaction dynamics and reasoning, while their lack of interpretability and limited generalization capabilities hinder their application in safety-critical scenarios.
Large Language Model for Decision-making. Large Language Models (LLMs), known for their human-like reasoning and commonsense generalization, have recently demonstrated potential beyond traditional natural language processing tasks  [16, 17]. In recent years, LLMs have been utilized as employed as end-to-end agents capable of autonomous decision-making. For cooperative driving tasks, Hu et al. proposed AgentsCoDriver [11], a LLM-driven lifelong learning framework to empower CAV decision-making. Meanwhile, CoDrivingLLM, proposed by Fang et al. [10], is designed to adapt different levels of Cooperative Driving Automation, which could effectively address the dilemma of cooperative scenarios. While LLMs have demonstrated remarkable potential, significant challenges remain. Their high computational demands and the complexity of real-time decision-making introduce latency issues, which limit their application in time-critical environments.
@@ III Problem Formulation
The cooperative decision-making problem for multiple CAVs can be modeled as a partially observable Markov decision process (POMDP) [21]. We define the POMDP using the tuple $\mathcal{M}_{\mathcal{G}}=(\mathcal{V},S,[\mathcal{O}_{i}],[\mathcal{A}_{i}],\mathcal{P},[r_{i}])$, where $\mathcal{V}$ represents the finite set of all controlled agents (CAVs), and $S$ denotes the state space encompassing all agents. $\mathcal{O}_{i}$ represents the observation space for each agent $i\in\mathcal{V}$, $\mathcal{A}_{i}$ denotes the action space, and $r_{i}$ is the reward associated with CAV $i$. The transition distribution is represented by $\mathcal{P}$. At any given time, each agent $i$ receives an individual observation $\mathcal{O}_{i}$ and selects an action $a_{i}\in\mathcal{A}_{i}$ based on a policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$.
III-A Observation Space
@@
where $x_{k}$ $y_{k}$ $v^{x}_{k}$ and $v^{y}_{k}$ are the longitudinal and lateral positions and speeds, respectively. $\phi_{k}$ is the vehicle’s heading angle. The overall observation space of the system is the combined observation of all CAVs, i.e., $\mathcal{O}=\mathcal{O}_{1}\times\mathcal{O}_{2}\times\cdots\times\mathcal{O}_{|\mathcal{V}|}$.
III-B Action Space
Given that the strength of LLMs lies in their reasoning capabilities based on world knowledge rather than numerical computation, we design the decision actions of CAVs as discrete semantic decisions rather than direct vehicle control actions. The action space $\mathcal{A}_{i}$ for agent $i$ is defined as a set of high-level control decisions, including $\{\textit{slow down},\textit{cruise},\textit{speed up},\textit{change left},\textit{change right}\}$. Once a high-level decision is selected, lower-level controllers generate the corresponding steering and throttle control signals to manage the CAVs’ movements. The overall action space is the combination of actions from all CAVs, i.e., $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{|\mathcal{V}|}$.
III-C Reward function
In this paper, with the objective of ensuring that all agents pass through the merging area safely and efficiently, based on previous work [18],the reward of $i$th agent at the time step $t$ is defined as follows:
@@ IV Method
In this section, the whole LDPD framework is first summarized. Then the teacher agent and student agent are introduced respectively.
@@ IV-A Framework Overview
The LDPD framework is designed to facilitate the learning and exploration process of multi-agent systems with distilled knowledge from LLM, especially in complex cooperative driving scenarios. As shown in Fig.2, the LDPD framework comprises two main components: the teacher agent and the student agents. The teacher agent facilitates efficient policy learning for multiple student agents through its action demonstrations. The student agents, composed of multiple small policy networks, form an agent group where each network is responsible for controlling an individual CAV. Throughout this process, the student agents efficiently distill and apply the knowledge from the LLM, enabling rapid learning and adaptation.
@@ IV-B1 Observation Enhancer
Additionally, based on the state information of the CAV, we predict and infer the driving intentions of each CAV, including the desired target lane and intended driving behavior. At time $t$, the driving intention information for CAV $i$ is represented as $In_{i,t}=\{lane^{\prime}_{i,t},behavior_{i,t}\}$, where $lane^{\prime}_{i,t}$ and $behavior_{i,t}$ refer to the desired lane intention and the desired behavior intention, respectively.
@@ IV-B2 Core Planner
We use the conflict checking function as an example to introduce our tools: After the LLM obtains the semantic description of the scenario, it will call the conflict checking tool conflict_check()conflict\_check() before initiating reasoning to detect potential safety risks.
@@
The conflict checking tool first retrieves all potential conflicts in the current environment and assesses the risk level based on the current states of the conflicting parties. Here, we utilize the time-to-conflict-point (TTCP) difference [10] as a surrogate metric for the risk level, which is widely used to evaluate collision risks at potential conflict points between vehicles, calculated as follows:
@@
A=LLM(Sce,conflict_info)A=LLM(Sce,conflict\_info) (6)
@@ IV-B3 Safety Checker
Firstly, the formula is designed to calculate the priority score of each CAV [18]:
@@
With the priority scores of CAVs calculated above, at time step $t$, the safety checker first produces a priority list ${P}_{t}$, which consists of a list of priority scores and their corresponding ego vehicles. Then, the safety checker sequentially examines the future trajectory of the vehicles in the list to determine if there are any conflicts between vehicles. We use the Intelligent Driver Model (IDM) [23] to predict the longitudinal acceleration of HVs, based on their current speed and distance headway, and we utilize the MOBIL lane change model [24] to predict the lateral behavior of HVs.
@@
where ${A}_{available}$ is a set of available actions at time step $k$ for the selected agent, and ${d}_{sm,k}$ is the safety margin at prediction time step $k$. The safety margin is defined as follows:
@@ V-C1 Overall Performance
The training curves for all methods under different conditions and scenarios are shown in Fig.5. As observed, in all experimental scenarios, our method (LEMAA2C) consistently outperforms the MARL baselines. Furthermore, they demonstrate significantly higher exploration efficiency in the early stages of training, which effectively boosts the training efficiency of the student agents. This strongly validates the effectiveness of our approach. Meanwhile, we also compare our method with some LLM-based approaches. As shown in Table I, our method exhibits the best performance across various tests, owing to its stronger exploration capability in the later stages of training.
@@
TABLE I: Performance comparison of different models under varying difficulty levels.
Scenario Model Easy Medium Hard
Scenario 1 GPT-3.5 [28] 91% 69% 23%
GPT-4o [29] 98% 82% 58%
DiLu [30] 96% 77% 60%
Ours 100% 96% 87%
Scenario 2 GPT-3.5 92% 72% 31%
GPT-4o 100% 78% 30%
DiLu 99% 82% 34%
Ours 100% 81% 67%
@@ V-C2 Safety Performance
The model’s overall performance is tested under various experimental conditions, as shown in Tab.II. Post Encroachment Time (PET), a key safety metric that quantifies the interaction and risk level between two vehicles, is used as one of the primary evaluation criteria. It is evident that our method consistently achieves the lowest collision rate across all tests. We also evaluate the safety performance of the teacher agent under different conditions and compare it with the student agents, as shown in Tab. III. In most scenarios, LAMAA2C outperforms LLM in the final performance, demonstrating that our method possesses strong exploration capabilities that further enhance performance. We believe this is due to the fact that as training progresses, the student model is able to integrate its own exploration and learning through policy updates more freely, in addition to receiving guidance from the teacher.
TABLE II: Performance across scenarios for various methods under different difficulty settings
@@
TABLE III: Performance of teacher and students under varying difficulty levels.
@@ V-C3 Efficiency Performance
The average speed of all CAVs, shown in Tab.II, is analyzed to assess the efficiency of cooperative driving across different methods. We observe that as the difficulty of tasks increases, the average speed of the CAVs decreases, which likely represents a compromise made to ensure safety over efficiency. However, compared to other baseline methods, such as MAPPO and MAACKTR, our approach not only maintains better safety performance but also ensures driving efficiency. This demonstrates that our method effectively balances driving speed with safety.
@@ V-C4 Cross-Validation
To evaluate the adaptability and generalization capabilities of the learned strategies, we test the models in task environments beyond their training conditions. For example, models trained under the hard mode are now tested on easy and medium mode tasks, with results shown in Tab.IV. We observe that models trained in hard mode not only outperform the original medium mode models in terms of safety and efficiency but also excel in other scenarios. This demonstrates the exceptional adaptability of our model.
TABLE IV: Performance comparison of models across different scenarios and difficulty levels.
Original Level Applied Level Evaluation Reward Collision Rate
Sce-1 Sce-2 Sce-1 Sce-2
HARD Simple 62.66 46.18 0.02 0.13
Medium 51.16 70.78 0.07 0.00
MEDIUM Simple 69.31 61.64 0.00 0.02
@@ VI Conclusions
To enhance exploration and learning efficiency in cooperative decision-making for CAVs, we introduced LDPD, a language-driven policy distillation framework to facilitate MARL agents learning. Within this framework, the language model, as the teacher, makes complex reasoning decisions based on observations from student agents and stores these decision-making experiences. Student agents learn the teacher’s prior knowledge through gradient updates in their policy networks. Extensive experiments demonstrate that distilled information from the teacher significantly enhances the exploration efficiency of student agents. Compared to baseline methods, there is a notable overall improvement in the agents’ performance, eventually surpassing the expert decision-making level of the teacher. In the future, we plan to extend our method to test additional MARL algorithms and apply it across a broader range of decision-making scenarios.